OpenStudy (blackops2_143):

Which fractions are equivalent to -4 - 9 ? Choose all answers that are correct. A. -2 - 5 B. -8 - 18 C. 16 - -36 D. 20 - 45

OpenStudy (blackops2_143):

I think maybe one of them is c?

pooja195 (pooja195):

-4/9 A. Can it be reduced to -2/5?

OpenStudy (blackops2_143):

ok is c right too

pooja195 (pooja195):

C is right

pooja195 (pooja195):

Find another one.

jimthompson5910 (jim_thompson5910):

If you want to know if two fractions are equivalent, here's what you do let's say we want to find out if 4/9 is equal to 2/5 step 1) set the two fractions equal to each other. At this point, we don't know if they are actually equal or not. This is what we want to test. This first step is basically making the claim "these two fractions are equal". The other steps will lead up to proving the claim true, or proving the claim false. So for now we have 4/9 = 2/5 step 2) cross multiply. So we go from 4/9 = 2/5 to 4*5 = 9*2 step 3) simplify each side that turns `4*5 = 9*2` into `20 = 18` which is FALSE Since the last equation is false, and all of the equations set up are equivalent, this makes the first equation false. So 4/9 is NOT equal to 2/5